OpenStudy (anonymous):

HELP (will give medal) 1) 4|2y-3|-1=11

OpenStudy (agent47):

4|2y-3|=11+1 4|2y-3|=12 |2y-3|=3

OpenStudy (agent47):

That breaks down into: (2y-3)=3 and -(2y-3)=3 2y = 6, y=3 -2y+3=3, -2y=0, y=0

OpenStudy (agent47):

So your answer is: y=0 or y=3